The work functions for several metals are shown in the table below. If 265 nm light is incident on each of these metals, determine: which metals exhibit the photoelectric effect.

Element

Chemical Symbol

Work Function (eV)

1

Barium

Ba

2.48

2

Beryllium

Be

3.92

3

Cesium

Cs

1.92

4

Gold

Au

4.82

5

Lithium

Li

2.35

6

Mercury

Hg

4.53

7

Nickel

Ni

5.01

Group of answer choices

All the metals listed will exhibit the photoelectric effect.

All the metals listed except nickel will exhibit the photoelectric effect.

All the metals listed except nickel and gold will exhibit the photoelectric effect.

Of the metals listed only cesium, lithium, barium and beryllium will exhibit the photoelectric effect.

Of the metals listed only cesium, lithium and barium will exhibit the photoelectric effect.

Of the metals listed only cesium and lithium will exhibit the photoelectric effect.

Of the metals listed only cesium will exhibit the photoelectric effect.
Answer #1

The work function is the minimum energy required to remove electron, i.e, to show photoelectric effect.

Now the incident light wavelength, \lambda=265nm which corresponds to energy, E=\frac{hc}{\lambda}=\frac{6.626*10^{-34}*3*10^{8}}{265*10^{-9}}J=75.01*10^{-20}J=4.69eV

So the incident light energy is less than the work function of Gold and Nickel.

So the right answer is: 3) All the metals listed except nickel and gold will exhibit the photoelectric effect.